APPLICATION OF LIE ALGEBROIDS ON FIELD THEORIES

JAIME CAMACARO

Abstract

We will show some examples of how the Lie algebroids can play an interesting role in the study of gauge fields theories. Our talk will be based on four examples: Yang–Mills theory, Topological sigma models, Poisson sigma models and open bosonic string, where the Lie algebroid structure can be easily recognized and used to obtain an algebra which take in consideration the gauge structure.

Jaime CAMACARO Departamento de Física Teórica, Facultad de Ciencias Universidad de Zaragoza SPAIN e-mail: jcama@wigner.unizar.es

Thanks: JRC acknowledges the financial support of the Agencia Española de Cooperación Internacional, under an AECI scholarship

Keywords: BV master equation, Gauge theories, Lie algebroids